1.a. Wheat is an important agricultural commodity, and the wheat market has been studied extensively by agricultural economists. From statistical studies, we know that for 1981 the supply curve for wheat was approximately as: Supply: QS = 1800 + 240P, where price is ineasured in nominal dollars per bushel and quantities in millions of bushels per year. These studies also indicate that in 1981, the demand curve for wheat was: Demand: QD = 3550 -266P. Determine the market-clearing price and quantity of wheat for 1981.

b. Use supply and demand curves to illustrate how each of the following events would affect the price of butter and the quantity of butter bought and sold: (i) an increase in the price of margarine; (ii) an increase in the price of milk; (iii) a decrease in average income levels.


1
Expert's answer
2021-03-08T09:22:24-0500

(a) The market equilibrium occurs when "Q_d=Q_s", then:


"3550-266P=1800+240P,""1750=506P,""P=\$3.46."

Therefore, the market-clearing price of the wheat is $3.46.

Finally, substituting "P" into "Q_d" we can find quantity of the wheat:


"Q=3550-266\cdot\$3.46=2630."


(b) (i)



Since butter and margarine are substitute goods, an increase in the price of margarine will lead to increase in the consumption of butter. As a result, the equilibrium price of butter will rise from "P_1" to "P_2", the equilibrium quantity of butter will rise from "Q_1" to "Q_2" and the demand curve for butter will shift to the right (from "D_1" to "D_2").

(ii)


An increase in the price of milk (the butter's main ingredient) will lead to increase the cost of producing of butter and, hence, will reduce the supply of butter. As a result, the equilibrium price of butter will rise from "P_1" to "P_2", the equilibrium quantity of butter will fall from "Q_1" to "Q_2" and the supply curve for butter will shift to the left (from "S_1" to "S_2").

(iii)


Since a butter is a normal good, the decrease in average income will lead to decrease in quantity of butter purchaised. As a result, the equilibrium price of butter will fall from "P_1" to "P_2", the equilibrium quantity of butter will fall from "Q_1" to "Q_2" and the demand curve for butter will shift to the left (from "D_1" to "D_2").